OpenStudy (anonymous):

I have 2.00 L of H2 at 475 torr, and 1.00L of N2 at 152 torr. If I combined the two into a 3.00L container, what would be my total pressure in torr? I'm looking for the process, not just the solution. Thanks for reading!

OpenStudy (anonymous):

ok let me try to crack this one fast: pV=nRT so n doesnt change cause there is no loss to surroundings: n= pV/RT index 1 before and index 2 after mixing n1=n2 p1V1/RT1=p2V2/RT2 you can cross out R and T cause they are constant so you got p1V1=p2V2 since you have 2 components: p1V1+p2V2=p3V3 475 torr * 2 L + 1 L * 152 torr = p3 * 3 L p3 = 367,33 torr that is if i got it rigt cause im a bit tired its 1:33 AM

OpenStudy (anonymous):

Doing the math on my end, but the process looks right. I'll get back with my solution. Thanks for the reply!

OpenStudy (anonymous):

he he he i also remembered easier way to come to same result :D 1/3 * 152 + 2/3 * 475 = 367,33 torr :D

OpenStudy (anonymous):

367.33 torr is what I got. Your second solution is much simpler! I think I can remember that. Thanks again Kryten!

OpenStudy (anonymous):

it comes from that famous law: p(partial) = p * x and here x is volume parts cause we are talking about gasses :D

OpenStudy (anonymous):

Would that be Dalton's Law of Partial Pressures?

OpenStudy (anonymous):

yup :)